ta-1754 e0001 A B C D
Betas Answer the questions below for assets A to D shown in the table: EEB a. What impact would a 11% increase in the market return be expected to have on each asset's return? b. What impact would a 8% decrease in the market return be expected to have on each asset's return? c. If you believed that the market return would increase in the near future, which asset would you prefer? d. If you believed that the market return would decrease in the near future, which asset would you prefer? a. If the market return increased by 11%, the impact to the return of asset A is negative for a decrease in the return.) %. (Round to one decimal place. Enter a positive percentage for an increase and a
Answer #1

(a) impact of increase in market return by 11%

Impact on stock's expected return formula = Beta*Impact on market return

Impact on A = -0.1*11%

-1.10%

impact on B = 0.7*11%

7.70%

Impact on C= -0.5*11%

-5.50%

Impact on D = 1.4*11%

15.40%

So A return decreased by 1.10%

B Return increased by 7.7%

C return decreased by 5.5%

D return increased by 15.4%

(b) impact of decrease in market return by 8%

Impact on A = -0.1* -8%

0.80%

impact on B = 0.7*-8%

-5.60%

Impact on C= -0.5*-8%

4.00%

Impact on D = 1.4*-8%

-11.20%

So A return increased by 0.80%

B Return decreased by 5.60%

C return increased by 4.0%

D return decreased by 11.20%

(c)

Beta denotes relationship between stock and market. it denotes ratio by which an asset return will be effected by change in market return

If beta is positive, asset return will move in positive direction. if beta is negative, asset return will move in opposite direction.

If market is expected to increase, we also want to increase in asset return. then security having positive beta will be selected.

So, Asset B and D would be preferred.

(d)

If market is expected to decline, then also we want to make our return more. so we will choose security with negative beta, so that our Asset return will move in opposite direction means positive return.

So we will choose Asset A and C